OpenStudy (anonymous):

Write the first derivative of the given function. f(x) = x^2[ln(x)]^14 I got f '(x) = 2x(lnx)^13 + (7+lnx) For which x-values does f(x) have horizontal tangents? (Enter your answers as a comma-separated list.) x =

OpenStudy (anonymous):

14 x (lnx)^13+2x(lnx)14

OpenStudy (anonymous):

yeah same thing, I just pulled out the common factor

OpenStudy (anonymous):

2x(lnx)^13(7+lnx)

OpenStudy (anonymous):

not +(7+lnx)

OpenStudy (anonymous):

Oh sorry, yeah I have that on my paper here I typed it wrong my real issue is finding where F'(x)=0

OpenStudy (anonymous):

not a problem 2x(lnx)^13(7+lnx)=0 7=-lnx 7=ln(1/x) e^7=1/x x=1/e^7

OpenStudy (anonymous):

I tried that but apparently it's wrong :(

zepdrix (zepdrix):

 Hmm do these work maybe? :O